# Contest for system observability as an infinitely repeated game

### v2

@@
Abstract
This paper studies a system security problem in the context of observability based on a two-person non-cooperative infinitely repeated game. Both the attacker and the defender have means to modify the dimension of the unobservable subspace, which is set as the value function. Utilizing tools from geometric control, we construct the best response sets considering one-step and two-step optimality respectively to maximize or minimize the value function. We find that the best response sets are not single-valued maps, resulting in a variety of game outcomes. In addition, we establish a necessary and sufficient condition for Nash equilibrium. The game analysis reveals that the defender plays a pivotal role in steering the game towards equilibrium, while the attacker can disrupt the equilibrium by sacrificing immediate gains.
keywords:
Observability, Linear system, Repeated games, Nash equilibrium, Geometric control
thanks: This work was supported in part by the National Natural Science Foundation of China with No. 62373245, in part by the National Key Research and Development Program of China with No. 2023YFB4706800, and in part by the Program of Shanghai Education Commission, China. Partial results of this paper were presented at IFAC World Congress 2023. Corresponding author Lin Wang.
@@ 1 Introduction
In recent years, more and more attention has been paid to the security of control systems. Remote sensors are vulnerable to attacks, which intend to deteriorate system performance by manipulating the transmitted data while remaining stealthiness [17]. Those attacks may have severe consequences. For example, in the cyber attack on the Ukrainian power grid in 2015, the attacker operated several of the circuit breakers in the grid and jammed the communication network to cause a large scale blackout and keep the operators unaware [21]. If attack can be found in time, the damage can be reduced.
@@
However, it should be emphasized that the above research focuses on whether the system can reconstruct the initial state qualitatively. They did not study dimension change of the unobservable subspace quantitatively, which can greatly impact the effectiveness of stealthy attacks. In this context increasing the dimension of the unobservable subspace can facilitate covert attacks, since the attack vector space for stealthiness is getting bigger [9], [26], [13]. In [9] and [26], subspace methods are used to construct undetectable attacks, yet without altering the dimension of the unobservable subspace. In [13], the attacker enhances the effect of stealthy attack by masking sensor signals to increase the dimension of the unobservable subspace.
In recent years, the game approach has shed new light on the system security problem [28], [10]. When the attacker and the defender have limited information, the game becomes partially observable. In [8], a partially observable stochastic game is studied and the authors give a representation of uncertainty encountered by the defender. In [27], an $\epsilon$-Stackelberg partially observable game model is built to prevent state information leakage. However, observability is considered as a game setting in the above research and few game research directly involves confrontation on the observability between attackers and defenders. In [12], a game approach is used to study the stealthy attack problem in which the attacker masked the sensors and the defender reinforced the sensors. However, in the utility of the players, observability is abstracted into different values without considering the relationship between observability and structural matrices of the system [12].
@@
Additionally, most existing game models use continuous value functions such as quadratic functions, and equilibrium solutions are obtained through methods such as dynamic programming in [23] and Q-learning in [18]. However, these methods are not suitable for discrete value functions, such as dimension of the unobservable subspace.
To address these limitations, an infinitely repeated game between the attacker and the defender is proposed, where the attacker aims to destroy system observability by maximizing the dimension of the unobservable subspace, while the defender tries to protect it. The main contributions are as follows.
(1) By taking the dimension of unobservable subspace as the value function, derivations and algorithms to maximize or minimize the dimension are given. The game antagonism based on observability is quantitatively studied.
(2) We investigate the impact of considering one-step and two-step optimality on game outcomes. We reveal that the best response sets are multi-valued, leading to diverse outcomes. Additionally, we give the condition for Nash equilibrium applicable to both scenarios.
(3) We analyze the disparities in the upper limits of capabilities between the attacker and defender. The defender dominates the game to achieve Nash equilibrium, while the attacker can forgo short-term benefits to prevent the establishment of the equilibrium, in order to gain long-term benefits.
Based on [24], this paper further studies the infinitely repeated game by incorporating two-step optimization in addition to one-step optimization. We find different action choices in the best response sets will lead to different game results. We analyze the Nash equilibrium condition and find the different effects of the attacker and defender on game equilibrium respectively.
The rest of paper is organized as follows. In Section 2, we formulate the system model and the game problem. In Section 3, derivations and algorithms are given to get the best response sets for both players. Section 4 analyzes the equilibrium and results for the game considering one-step optimality and two-step optimality respectively. In Section 5, examples are given to illustrate various game results. Section 6 is a brief conclusion. A summary of notations is provided in Table 1.
Table 1: Notations
@@ 2.1 Modeling of the system
With the relative degree, it is well known that after a coordinate change $x\rightarrow\left(\begin{array}[]{l}z\\ \xi\end{array}\right)$, where $\xi=\left[\xi_{1}^{1},\xi_{2}^{1},\cdots,\xi_{r_{1}}^{1},\cdots,\xi_{1}^{m},\xi_{2}^{m},\cdots,\xi_{r_{m}}^{m}\right]^{\top}$, system (1) can be rewritten as the normal form
$\dot{z}$ $=Nz+E\xi+B_{a}^{\prime}u_{a},$ (2)
@@
where $K_{d}\in\mathbb{R}^{m\times s}$, $U_{d}\in\mathbb{R}^{m\times(n_{0}-s)}$ are determined by the defender; $K_{a}\in\mathbb{R}^{k\times s}$, $U_{a}\in\mathbb{R}^{k\times(n_{0}-s)}$ are determined by the attacker; $u_{0}$ is the input that maintains the normal operation of the system and is controlled by a basic controller, which is illustrated in Fig. 1(a). Then
@@
Figure 1: Comparison of the model in the normal form and in the observability equivalent system.
@@
$R+LU_{d}\triangleq{H}.$
Since $L$ is a non-singular matrix according to the definition of relative degree, $\forall{H}\in\mathbb{R}^{m\times n}$, there exists $U_{d}=L^{-1}\left({H}-R\right)$. Thus the defender can completely control ${H}$. If the system is observable, when $y\equiv\mathbf{0}$, there is $\left(\begin{array}[]{l}z\\ \xi\end{array}\right)\equiv\mathbf{0}$. Consider $y\equiv\mathbf{0}$, according to the coupling relationship between $y_{i}$ and $(\xi_{1}^{1},\cdots,\xi_{r_{1}}^{1},\cdots,\xi_{1}^{m},\cdots,\xi_{r_{m}}^{m})=\xi^{\top}$, we have $\xi\equiv\mathbf{0}$. Thus we only need to prove $z\equiv\mathbf{0}$, where $z$ satisfies
@@
Define $Hz=\hat{y}$ and $n_{0}-s=n$. The condition for completely observable becomes: if $\hat{y}\equiv\mathbf{0}$, there is $z\equiv\mathbf{0}$, which is also the condition for the following system to be observable:
$\begin{split}\dot{z}&=Nz+B_{a}^{\prime}U_{a}z,\\ \hat{y}&=Hz,\end{split}$ (9)
where $z\in\mathbb{R}^{n}$ and $\hat{y}\in\mathbb{R}^{m}$ are state and output of the observability equivalent system respectively; $N\in\mathbb{R}^{n\times n}$ is system matrix and $B_{a}^{\prime}\in\mathbb{R}^{n\times k}$ is input matrix. Now the problem becomes that the attacker wants to damage system observability using state feedback control $U_{a}$, while the defender wants to protect system observability by controlling $H$, which is illustrated in Fig. 1(b).
@@ 2.2 Game formulation
Define an infinitely repeated game as a tuple $(N,A,J)$. $N=\{attacker(a),defender(d)\}$ is the set of players. $A=A^{a}\oplus A^{d}$ is the action set of players, where $A^{a}=\{U_{a}\in\mathbb{R}^{k\times n}\},A^{d}=\{H\in\mathbb{R}^{m\times n}\}$. $J=\{J^{a},J^{d}\}$ is the utility function of the attacker and the defender. Define the utility function of the attacker in epoch $i$ as
$J^{a}(U_{ai},H_{i})$ $=\operatorname{dim\,Ker}\!\!\left[\begin{array}[]{c}H_{i}\\ H_{i}\left(N+B_{a}^{\prime}U_{ai}\right)\\ \vdots\\ H_{i}\left(N+B_{a}^{\prime}U_{ai}\right)^{n-1}\end{array}\right]$
$\overset{\triangle}{=}\operatorname{dim\,Ker}\Omega(U_{ai},H_{i}),$
where $\Omega$ is the observability matrix of system (9), $U_{ai}$ and $H_{i}$ are actions of players in epoch $i$. As for the defender,
$J^{d}(U_{ai},H_{i})$ $=-J^{a}(U_{ai},H_{i}).$ (10)
Define the value function $\Phi$ in epoch $i$ as the dimension of unobservable subspace
$\Phi(U_{ai},H_{i})=\operatorname{dim\,Ker}\Omega(U_{ai},H_{i}).$ (11)
Thus the attacker aims to maximize the value function, while the defender seeks to minimize. Two players update actions asynchronously and an epoch is defined once a player acts. Repetitive games can effectively reflect the interaction between the attacker and defender, guiding both sides in designing strategies to gain an advantage in this tug-of-war relationship.
@@ 3.1 The best response set of the attacker
Denote $\mathcal{V}\triangleq\operatorname{Ker}\,\Omega$. It is apparent that $\mathcal{V}$ is a controlled invariant subspace contained in $\operatorname{Ker}H$. Among all controlled invariant subspaces contained in $\operatorname{Ker}H$, there is a maximal one denoted as $\mathcal{V}^{*}$, thus,
$\max_{U_{a}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)=\operatorname{dim}\mathcal{V}^{*}.$ (13)
A lemma is given to find $\mathcal{V}^{*}$.
Lemma 1.
@@
By classical results on maximal controlled invariant subspace, we can conclude that $U_{a}\in\mathbb{R}^{k\times n}$ maximizes the dimension of unobservable subspace if and only if it is a friend matrix of $\mathcal{V}^{*}$, i.e.,
$U_{a}^{*}\in\mathcal{F}(\mathcal{V}^{*}),$ (15)
where $\mathcal{F}(\mathcal{V}^{*})$ is defined in Definition 1. In Algorithm 1, a way using pseudo inverse to find a friend matrix is given. For most cases, the friend matrix is not unique. Since in our case, all friend matrices are equivalent, so we use the pseudo inverse to get the explicit form for $U_{a}^{*}$. The calculation steps of $U_{a}^{*}$ is summarized in Algorithm 1.
Algorithm 1 Algorithm for the attacker to maximize the unobservable subspace
  Set $\operatorname{Im}V_{0}=\operatorname{Ker}H$. Calculate $\operatorname{Im}Z_{i}=\left(\operatorname{Ker}\left[\begin{array}[]{c}V_{0}^{\top}\\ B_{a}^{\prime\top}\end{array}\right]\right)^{\top}$ Calculate $\operatorname{Im}V_{1}=\operatorname{Ker}\left[\begin{array}[]{c}H\\ Z_{i}N\end{array}\right]$.
  while $\operatorname{Im}V_{i}\neq\operatorname{Im}V_{i+1}$ do
     Set $V_{i}=V_{i+1}$. Calculate $\operatorname{Im}Z_{i}=\left(\operatorname{Ker}\left[\begin{array}[]{c}V_{i}^{\top}\\ B_{a}^{\prime\top}\end{array}\right]\right)^{\top}$. Calculate $\operatorname{Im}V_{i+1}=\operatorname{Ker}\left[\begin{array}[]{c}H\\ Z_{i}N\end{array}\right]$.
  end whileSet $V=V_{i}=\left[\begin{array}[]{llll}v_{1}&v_{2}&\ldots&v_{r}\end{array}\right]$.Calculate $[X;U]=\text{pinv}\left(V\quad B_{a}^{\prime}\right)NV$, choose the last $\quad(n-r)$ columns as $U$.Calculate $U_{a}=-U\text{pinv}(V)$.

### v3

@@
Contest for system observability as an infinitely repeated game
XU Yueyue $\cdot$ ZHOU Panpan $\cdot$ WANG Lin $\cdot$ LIU Zhixin $\cdot$ HU XiaomingXU Yueyue
Department of Automation, Shanghai Jiao Tong University, Shanghai 200240, China; Department of Mathematics, KTH Royal Institute of Technology, Stockholm 10044, Sweden. Email: merryspread99@sjtu.edu.cn.
WANG Lin (Corresponding author)
Department of Automation, Shanghai Jiao Tong University, Shanghai 200240, China. Email: wanglin@sjtu.edu.cn.
ZHOU Panpan $\cdot$ HU Xiaoming
Department of Mathematics, KTH Royal Institute of Technology, Stockholm 10044, Sweden. Emails: panpanz@kth.se; hu@kth.se.
LIU Zhixin
Key Laboratory of Systems and Control, Academy of Mathematics and Systems Science, CAS, Beijing 100190, China. Email: lzx@amss.ac.cn. This work was supported in part by the National Natural Science Foundation of China with No. 62373245, in part by the National Key Research and Development Program of China with No. 2023YFB4706800, and in part by the Dawn Program of Shanghai Education Commission, China. Partial results of this paper were presented at IFAC World Congress 2023.
DOI:
Received: x x 2025 / Revised: x x 2025
[…]
This paper studies a system security problem in the context of observability based on a two-person noncooperative infinitely repeated game. Both the attacker and the defender have means to modify the dimension of the unobservable subspace, which is set as the value function. Utilizing tools from geometric control, we construct the best response sets considering one-step and two-step optimality respectively to maximize or minimize the value function. We establish a unified necessary-and-sufficient condition for Nash equilibrium that holds for both one-step and two-step optimizations. Our analysis further uncovers two evolutionary patterns, lock and loop modes, and shows an asymmetry between defense and attack. The defender can lock the game into equilibrium, whereas the attacker can disrupt it by sacrificing short-term utility for longer-term advantage. Six representative numerical examples corroborate the theoretical results and highlight the complexity of possible game outcomes.
[…]
Observability, linear system, repeated games, nash equilibrium, geometric control.
@@ 1 Introduction
In recent years, more and more attention has been paid to the security of control systems. Remote sensors are vulnerable to attacks, which intend to deteriorate system performance by manipulating the transmitted data while remaining stealthy [1]. Those attacks may have severe consequences. For example, in the cyber attack on the Ukrainian power grid in 2015, the attacker operated several of the circuit breakers in the grid and jammed the communication network to cause a large scale blackout and keep the operators unaware [2]. If the attack can be detected in time, the damage can be reduced.
@@
However, it should be emphasized that the above research focuses on whether the system can reconstruct the initial state qualitatively. They did not study dimension change of the unobservable subspace quantitatively, which can greatly impact the effectiveness of stealthy attacks. In this context increasing the dimension of the unobservable subspace can facilitate covert attacks, since the stealthy attack vector space is getting larger [12], [13], [14]. In [12] and [13], subspace methods are used to construct undetectable attacks, yet without altering the dimension of the unobservable subspace. In [14], the attacker enhances the effect of stealthy attack by masking sensor signals to increase the dimension of the unobservable subspace.
In recent years, the game approach has shed new light on the system security problem [15], [16]. When the attacker and the defender have limited information, the game becomes partially observable. In [17], a partially observable stochastic game is studied and the authors give a representation of uncertainty encountered by the defender. In [18], an $\varepsilon$-Stackelberg partially observable game model is built to prevent state information leakage. However, observability is considered as a game setting in the above research and few game research directly involves confrontation on the observability between attackers and defenders. In [19], a game approach is used to study the stealthy attack problem in which the attacker masked the sensors and the defender reinforced the sensors. However, in the utility of the players, observability is abstracted into different values without considering the relationship between observability and structural matrices of the system [19].
@@
Additionally, most existing game models use continuous value functions such as quadratic functions, and equilibrium solutions are obtained through methods such as dynamic programming in [23] and Q-learning in [24]. However, these methods are not suitable for discrete value functions, such as the dimension of the unobservable subspace.
To overcome the above limitations, this study models the confrontation between the attacker and the defender as an infinitely repeated game. The attacker attempts to undermine system observability by maximizing the dimension of the unobservable subspace, whereas the defender seeks to minimize that dimension so as to preserve observability. The main contributions are summarized below.
(1) We introduce the dimension of the unobservable subspace as the value function and develop closed-form best response algorithms for both players. By combining geometric control computations of controlled invariant subspaces with game model, this work quantifies the adversarial contest over system observability. In contrast to [25], which addresses only one-step optimality, our framework establishes an extended-horizon optimization formulation and derives analytical expressions.
(2) We establish a unified necessary-and-sufficient condition for Nash equilibrium that holds for both one-step and two-step optimizations, expressed as a concise equality test. Because two players move in alternation, extending the optimization horizon beyond two steps provides no additional strategic information. Previous studies have not provided such a unified equilibrium criterion.
(3) We discover the defense–attack asymmetry and its resulting game-outcome patterns. The theoretical and experimental analysis shows that the defender can lock the game into equilibrium, whereas the attacker can disrupt it by postponing immediate gains. Two evolutionary paths, lock mode and loop mode, have been identified, providing new theoretical guidance for designing practical security counter-strategies. This finding highlights a fundamental asymmetry neglected in earlier work that usually assumes symmetric influence.
The rest of paper is organized as follows. In Section 2, we derive an observability equivalent system and formulate the game problem. In Section 3, derivations and algorithms are given to get the best response sets for both players. In Section 4, we derive the Nash equilibrium under both one-step and two-step optimality, then refine the game outcome modes and analyze the equilibrium characterization. In Section 5, examples are given to illustrate various game results. Section 6 is a brief conclusion. A summary of notations is provided in Table 1.
Table 1   Notations
@@ 2.1 Modeling of the system
With the relative degree, it is well known that after a coordinate change $x\rightarrow\left[\begin{array}[]{l}z\\ \xi\end{array}\right]$, where $\xi=\left[\xi_{1}^{1},\xi_{2}^{1},\cdots,\xi_{r_{1}}^{1},\cdots,\xi_{1}^{m},\xi_{2}^{m},\cdots,\xi_{r_{m}}^{m}\right]^{\top}$, system (1) can be rewritten as the normal form
$\dot{z}$ $=Nz+E\xi+B_{a}^{\prime}u_{a},$
@@
where $K_{d}\in\mathbb{R}^{m\times s}$, $U_{d}\in\mathbb{R}^{m\times(n_{0}-s)}$ are determined by the defender; $K_{a}\in\mathbb{R}^{k\times s}$, $U_{a}\in\mathbb{R}^{k\times(n_{0}-s)}$ are determined by the attacker; $u_{0}$ is the input that maintains the normal operation of the system and is controlled by a basic controller, which is illustrated in Figure 1(a).
@@
Figure 1 Comparison of the model in the normal form (a) and in the observability equivalent system (b).
Then we have
@@
$R+LU_{d}\triangleq{H}.$ (5)
Since $L$ is a non-singular matrix according to the definition of relative degree, $\forall{H}\in\mathbb{R}^{m\times n}$, there exists $U_{d}=L^{-1}\left({H}-R\right)$. Thus the defender can completely control ${H}$. If the system is observable, when $y\equiv\mathbf{0}$, there is $\left[\begin{array}[]{l}z\\ \xi\end{array}\right]\equiv\mathbf{0}$. Consider $y\equiv\mathbf{0}$, according to the coupling relationship between $y_{i}$ and $(\xi_{1}^{1},\cdots,\xi_{r_{1}}^{1},\cdots,\xi_{1}^{m},\cdots,\xi_{r_{m}}^{m})=\xi^{\top}$, we have $\xi\equiv\mathbf{0}$. Thus we only need to prove $z\equiv\mathbf{0}$, where $z$ satisfies
@@
Define $Hz=\widehat{y}$ and $n_{0}-s=n$. The condition for completely observable becomes: if $\widehat{y}\equiv\mathbf{0}$, there is $z\equiv\mathbf{0}$, which is also the condition for the following system to be observable:
$\begin{split}\dot{z}&=Nz+B_{a}^{\prime}U_{a}z,\\ \widehat{y}&=Hz,\end{split}$ (7)
where $z\in\mathbb{R}^{n}$ and $\widehat{y}\in\mathbb{R}^{m}$ are state and output of the observability equivalent system respectively; $N\in\mathbb{R}^{n\times n}$ is system matrix and $B_{a}^{\prime}\in\mathbb{R}^{n\times k}$ is input matrix. Now the problem becomes that the attacker wants to damage system observability using state feedback control $U_{a}$, while the defender wants to protect system observability by controlling $H$, which is illustrated in Figure 1(b).
@@ 2.2 Game formulation
Define an infinitely repeated game as a tuple $(N,A,J)$. $N=\{a,\,d\}$ is the set of players, where $a,\,d$ represent the attacker and the defender respectively. $A=A^{a}\oplus A^{d}$ is the action set of players, where $A^{a}=\{U_{a}\in\mathbb{R}^{k\times n}\},A^{d}=\{H\in\mathbb{R}^{m\times n}\}$. $J=\{J^{a},J^{d}\}$ is the utility function set of the attacker and the defender. Define the utility function of the attacker in epoch $i$ as
$J^{a}(U_{a,i},H_{i})=\operatorname{dim\,Ker}\!\!\left[\begin{array}[]{c}H_{i}\\ H_{i}\left(N+B_{a}^{\prime}U_{a,i}\right)\\ \vdots\\ H_{i}\left(N+B_{a}^{\prime}U_{a,i}\right)^{n-1}\end{array}\right]\overset{\triangle}{=}\operatorname{dim\,Ker}\Omega(U_{a,i},H_{i}),$
where $\Omega$ is the observability matrix of system (7), $U_{a,i}$ and $H_{i}$ are actions of players in epoch $i$. As for the defender, $J^{d}(U_{a,i},H_{i})=-J^{a}(U_{a,i},H_{i})$. Define the value function $\Phi$ in epoch $i$ as the dimension of unobservable subspace, i.e.,
$\Phi(U_{a,i},H_{i})=\operatorname{dim\,Ker}\Omega(U_{a,i},H_{i}).$ (8)
Therefore, the defender aims to minimize the value function, while the attacker seeks to maximize it. Two players update actions asynchronously and an epoch is defined once a player acts. This asynchronous decision making between the defender and the attacker reflects industrial-control practice: the defender first configures the observations; only once data start flowing can the attacker tamper with them, naturally producing a defense-then-attack cycle. Repetitive games can effectively reflect the interaction between the defender and attacker, guiding both sides in designing strategies to gain an advantage in this tug-of-war relationship.
Remark 2.2.
The usual discounted-sum criterion for repeated games is not adopted here, because the stage utility is the discrete dimension of the unobservable subspace, which is a quantity that is inherently non-additive. Instead, the analysis focuses on the unobservable dimension that persists once the play converges (or settles into a limit cycle), rather than on any discounted accumulation of these dimensions.
@@ 3.1 The best response set of the attacker
Denote $\mathcal{V}\triangleq\operatorname{Ker}\,\Omega$. Then $\mathcal{V}$ is a controlled invariant subspace contained in $\operatorname{Ker}H$. Among all controlled invariant subspaces contained in $\operatorname{Ker}H$, there is a maximal one denoted as $\mathcal{V}^{*}$, thus
$\max_{U_{a}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)=\operatorname{dim}\mathcal{V}^{*}(H).$ (10)
Here follows a lemma to find $\mathcal{V}^{*}$.
Lemma 3.1.
@@
Classical results on maximal controlled invariant subspaces show that $U_{a}\in\mathbb{R}^{k\times n}$ maximizes the dimension of the unobservable subspace iff it is a friend matrix of $\mathcal{V}^{*}$; formally,
$U_{a}^{*}\in\mathcal{F}(\mathcal{V}^{*}(H)),$
where $\mathcal{F}(\mathcal{V}^{*})$ is defined in Definition 2.1. Algorithm 1 constructs such a friend matrix via the pseudo inverse. In most cases, the friend matrix is not unique; hence the pseudo inverse provides a convenient explicit realization of $U_{a}^{*}$. The complete computational steps are summarized in Algorithm 1.
Algorithm 1 Attacker: maximization of the unobservable subspace
  Input system matrices $N$, $B_{a}^{\prime}$; the defender action $H$
  Set $\operatorname{Im}V_{0}=\operatorname{Ker}H.$ Calculate $\operatorname{Im}Z_{1}=\left(\operatorname{Ker}\left[\begin{array}[]{c}V_{0}^{\top}\\ B_{a}^{\prime\top}\end{array}\right]\right)^{\top}$, $\operatorname{Im}V_{1}=\operatorname{Ker}\left[\begin{array}[]{c}H\\ Z_{1}N\end{array}\right]$.
  Set $i=1$
  while $\operatorname{Im}V_{i}\neq\operatorname{Im}V_{i-1}$ do
     Set $i=i+1$. Calculate $\operatorname{Im}Z_{i}=\left(\operatorname{Ker}\left[\begin{array}[]{c}V_{i-1}^{\top}\\ B_{a}^{\prime\top}\end{array}\right]\right)^{\top}$, $\operatorname{Im}V_{i}=\operatorname{Ker}\left[\begin{array}[]{c}H\\ Z_{i}N\end{array}\right]$.
  end whileSet $V=V_{i}$.Calculate $[X;U]=\text{pinv}\left(V\,\,\,B_{a}^{\prime}\right)NV$, choose the last $(n-r)$ columns as $U$.Calculate $U_{a}^{*}=-U\text{pinv}(V)$.return $U_{a}^{*}$